# Evolutionary Optimization of Physics-Informed Neural Networks: Survey and Prospects

### v2

@@ III-A3 Sensitivity to Sample Size and Location
The training samples of the PINN model are typically uniformly drawn from the input domain and are equally important everywhere due to their physical correspondence with the spatio-temporal extent of the problem. While the use of mini-batch samples is a standard procedure in gradient-based training, it remains an open question how one can efficiently reduce the batch size (and hence the training cost) while maintaining the desired PINN model accuracy [91]. Using a large batch size is crucial for the accuracy of PINNs due to the potential need to learn high frequencies or complex structures in the a priori unknown target function. However, this inevitably increases training costs. Furthermore, oversampling on a high loss region may adversely slow convergence since it can be harder to learn such local complex gradient behaviour—a cause of stiffness during training [51, 52, 37, 38]. On the other hand, an overly small batch size will result in noisy (big jump in loss values) and inconsistent local gradient signals [92], thereby causing issues during gradient descent.
@@ III-B Challenges in Learning Generalizable PINN Model
With the simulation (hypothesis evaluation) and inference capability, PINNs are versatile and can be applied to various applications common in science such as model-predictive-control, scenario planning, and design exploration [93, 94, 95, 96]. The commonality to these applications is the need for multiple accurate and physics-compliant predictions across a series of related problems. These repeated evaluations usually differ by modifications to the physics scenario such as changes in geometric shape, PDE parameter, IC and BC.
@@ III-B1 How Generalizable Is A PINN Model?
Most parameterized PINN models (ref. to Subsection II-C; they are designed to learn from a parametric family of physics scenarios) make predictions on new scenarios unseen during training through interpolation. Contrary to its promise, the interpolation is usually physics-agnostic, with similar negative implications for generalizability as the predictions from typical DNNs [50]. Given the difficulties in optimizing a PINN model for a single physics scenario, the cost of performing physics-informed learning across a set of scenarios can be prohibitive [97]. The effectiveness of these parameterized PINNs for prediction of new scenarios then depends on the complexity of the problem. Moreover, the problem’s parameterization needs to be determined at the learning stage, which reduces the flexibility of the parameterized PINN model when applied to new physics scenarios.
@@ III-B2 How to Train A Generalizable PINN Model in The Data-Scarce Regime?
IV Physics-Informed Learning with Neuroevolutionary Algorithms
@@
Despite their theoretical advantage in diverse scientific domains, Section III highlighted several practical challenges in today’s PINN learning and optimization. In this section, we postulate that many algorithmic ideas developed in the computational intelligence community may be uniquely well suited for addressing issues plaguing PINNs. One exciting research direction is to synergise physics-informed neuroevolution with local gradient descent or least-squares signals, thereby leveraging the best of both worlds to quickly arrive at physically accurate and generalizable PINN models. In what follows, we delve deeper into this and other algorithmic pathways deemed worthy of exploration.
IV-A Exploration-Oriented Evolutionary Algorithms with Global Search Capacity
IV-A1 Neuroevolution
@@
While examining PINN optimization challenge, it becomes apparent that there is a need for new algorithms that can overcome issues such as local minima traps and vanishing gradients faced by conventional gradient descent algorithms, including the widely used SGD. Gradient descent is a point-based optimization method that is highly sensitive to the issues of local minima or learning bias magnified by PINN loss functions. In contrast, in silico evolutionary computation is a class of highly parallelizable approaches for guiding populations of candidate solutions towards near-optimal regions of search / parameter spaces via mechanisms inspired by biological evolution [108]. We shall broadly refer to any such algorithmic realization as an evolutionary algorithm (EA) and its application to the training of neural networks as neuroevolution.
@@
The operators of EAs are typically based on principles of natural selection. They do not require derivative computations of the loss functions being optimized and are therefore gradient-free. Importantly, as population-based methods, EAs offer global search capability with lesser tendency to get trapped in a single local minimum [109, 47]. What’s more, since their search is typically invariant to order-preserving transformations of the loss function [110], EAs are less sensitive to small (vanishing) gradients. For these reasons, they are deemed as potential alternatives to SGD, particularly in the context of PINN where the goal is to converge to physically precise solutions with close to zero loss [46]. In addition to its global search capacity, an evolved population naturally provides an ensemble of neural networks that affords the possibility of robust predictions with quantified uncertainty [111, 112].
In neuroevolution, EAs are often tasked with directly optimizing the parameters of neural networks. However, the large-scale nature of neural network parameters poses significant challenges, as EAs typically struggle to perform effectively under such conditions. To address this, several advanced techniques, such as sparse optimization [113] and coevolution [114], have been developed to enhance the performance of EAs for large-scale variable optimization. They provide valuable insights and inspiration for addressing large-scale optimization challenges within the context of PINE.
IV-A2 Pareto Optimization
@@
The existence of multiple loss terms—see Eqs. (2) and (3)—makes PINN training intrinsically a multi-objective optimization problem. Since the terms $\mathcal{L}_{data},\mathcal{L}_{pde},\mathcal{L}_{ic}$ and $\mathcal{L}_{bc}$ represent different physical quantities, their magnitudes may be incommensurable. In this case, the role of the scalarizing weights is not only to express preferences between the losses, but also to compensate for differences in their magnitudes. The possible conflict between the losses and the non-convexity of the aggregated physics-informed learning objective implies that a neural network, despite being a universal function approximator, is not guaranteed to find a solution that simultaneously minimizes all terms [44]. Moreover, the shape of the Pareto front formed by the multiple objectives is not known beforehand [115], raising concerns that a simple weighted-sum aggregation approach may not support all parts of the front [116].
@@
IV-A3 Memetic Algorithms
There is a clear theoretical rationale supporting physics-informed neuroevolution, yet there are only a handful of related works in the literature that explore such ideas. A barrier to the widespread use of EAs has been their high sample complexity, a consequence of updating and evaluating populations of solutions over multiple iterations. It is useful to consider hybridizing global evolutionary search with local gradient signals—thus inheriting the merits of both approaches. While the former can serve as an efficient basin-hopper for exploring rugged PINN landscapes, the latter facilitates rapid convergence to stationary points by exploiting local derivative information of loss functions.
@@ IV-A3 Memetic Algorithms
The crafting of synergistic optimization frameworks to meld gradient-based and gradient-free learning poses fundamental challenges, including determining interaction mechanisms that best harness the strengths of the two approaches. Key questions to address include balancing the intensity of SGD versus neuroevolution, ascertaining when gradient descent should take over from evolution, or selecting an appropriate subset of the evolving population for fine-tuning with gradient descent (as doing so for the entire population would be too expensive). Notably, the subfield of memetic computation in computational intelligence—where ‘memes’ are loosely defined as computationally-encoded units of information (e.g., local gradient signals) combined with global evolutionary search for effective problem-solving—offers a wealth of theories and methodologies that directly tackle these issues [129, 130].
IV-A4 Hardware-Accelerated Neuroevolution
@@
IV-B Neural Architecture and Hyperparameter Search
The choice of architectures and hyperparameters for PINNs has often been guided by experiential priors drawn from standard deep learning. However, a growing body of literature shows that accurate physics-informed learning may call for bespoke architectures. For example, contrary to popular belief, it was found that more hidden layers do not necessarily mean better performance and can sometimes be detrimental. An automated neural architecture search via a bilevel programming formulation revealed that, for certain governing differential equations, a shallow neural network with many neurons was more suitable for PINNs [133]. Similarly intriguing insights regarding the choice of activation function have also been uncovered. While squashing functions (e.g., tanh and sigmoid) and variants of the rectified linear unit are the mainstay in the deep learning literature, the use of Fourier features (i.e., sine and cosine) has gained prominence in physics tasks [84, 85], specifically for their ability to overcome spectral bias and capture periodic or high frequency patterns (e.g., shock fronts) common in various physics applications [134, 135, 136].
@@ IV-B Neural Architecture and Hyperparameter Search
Figure 6: The projectiles under the effect of Earth, Mars, and Moon are respectively solved by transfer neuroevolution—dubbed tNES. The convergence trends from SGD, xNES, and tNES optimization runs for selected scenarios are displayed (the bold lines indicate the mean convergence path of multiple independent runs and the shaded areas indicate their 10th-90th interpercentile ranges). Individual solutions given by multiple SGD and tNES (transfer from a relevant prior, i.e., source distribution obtained by solving a different gravity $g$ or launch angle $a_{0}$) optimization runs for one particular scenario ($g$=1.6, $a_{0}$=45) are plotted and compared against the ground truth.
IV-C Transfer Learning and Transfer Evolutionary Optimization
IV-C1 Learning PINN on Related Tasks
@@ IV-C2 Adaptive Transfer
IV-D Meta-Learning by Neuroevolution
IV-D1 Meta-Learning in PINN
Meta-learning algorithms provide a systematic framework for reusing knowledge from a distribution of learning tasks. This allows a meta-learned PINN model to quickly adapt to new physics-informed learning tasks with minimal training. Crucially, meta-learning offers a pathway towards a generalizable PINN over a broad class of physics scenarios, including different PDE families and geometries, as highlighted in Subsection III-B2. For example, the MAML and Reptile algorithms have been recently explored in the context of PINN [102, 103, 101]. They seek to meta-learn an optimal model initialization in the space of network weights from which the network can rapidly converge to a solution that minimizes the physics-informed loss of new tasks. This involves a task-specific fine-tuning procedure during the meta-learning stage through an inner-outer loop training framework, i.e., a bi-level optimization problem. The network weights are trained explicitly (outer loop) by optimizing the model performance over a batch of physics-informed learning tasks after one or a few gradient descent updates from the learned initialization (inner loop).
@@ IV-D2 Neural Baldwinism
Figure 8: Schematic to illustrate new tasks arising from family of PDEs problem: S1 change to new PDE ($\gamma$ and $\delta$ parameters), source term $q$, and IC profile for $t\in[0,2]$ (same time domain as train tasks), and S2 projection to longer time domain $t\in[0,4]$. The solution for new PDE tasks (they are visually indistinguishable from the ground truth) can be quickly obtained in the order of milli-seconds by the Baldwinian-PINN [50].
@@
Recently, a meta-learning PINN framework based on neuroevolution—dubbed Baldwinian PINN [50]—has been proposed to address the training challenges faced in conventional gradient-based meta-learning algorithms such as MAML. The Baldwinian-PINN provides a principled approach for optimizing a generic model initialization for network weights and other learning hyperparameters essential for achieving optimal performance across a distribution of physics-informed learning tasks. Interestingly, the Baldwinian-PINN limits task-specific learning to the output layer only [165, 166, 167], allowing it to solve new linear PDE problems with a 1-step pseudoinverse operation (or a few pseudoinverse steps for nonlinear PDE problems). The pseudoinverse operation is fast and exact. Unlike gradient-based MAML algorithms, the Baldwinian-PINN uses a global search strategy for neuroevolution, which can better handle the bi-level optimization problem with sparse and noisy rewards. It does not require unrolling and differentiating through the computation graph, making Baldwinian-PINN a more efficient strategy that scales well with the number of GPUs available. Fig. 7 illustrates the contrast between MAML and Baldwinian evolution for PINN [^2] Python source codes for Baldwinian-PINN are available at https://github.com/chiuph/Baldwinian-PINN..
@@
The Baldwinian-PINN can be re-used to effectively solve new physics scenarios within a diverse task distribution with orders of magnitude improvement in speed and accuracy compared to single PINN model. The effective generalization of a Baldwinian-PINN to an entire PDE family with diverse output patterns is demonstrated in Fig. 8, on 2 task scenarios: S1 shows successful learning of solution for unseen set of PDEs, which include changes to PDE parameters, source term, and IC profile; while S2 shows effective extrapolation of solution to a longer time domain. Interestingly, the Baldwinian-PINN maintain good accuracy on tasks from S2, whereby the model first learns the PDE solution for $t\in[0,2]$, before using the learned solution $u(x,t=2)$ as new IC for solving $t\in[2,4]$. This is achievable because Baldwinian evolution does not require parameterization for the tasks, and evolved Baldwinian-PINN can generalize to new ICs, BCs, and PDE source terms. This is tricky for existing parameterized PINNs as they require interpolation across a potentially infinitely large distribution of tasks (e.g. possible BCs or ICs), further emphasizing the merits of Baldwinian evolution for physics. These encouraging results suggest that similar approaches in the spirit of Baldwinian-PINN are worthy of further consideration, such as for the continual modelling of dynamical systems through their versatility in handling ICs and ability to rapidly model and stitch time windows together with minimal error.

### v3

@@ III-A3 Sensitivity to Sample Size and Location
The training samples of the PINN model are typically uniformly drawn from the input domain and are equally important everywhere due to their physical correspondence with the spatio-temporal extent of the problem. While the use of mini-batch samples is a standard procedure in gradient-based training, it remains an open question how one can efficiently reduce the batch size (and hence the training cost) while maintaining the desired PINN model accuracy [98]. Using a large batch size is crucial for the accuracy of PINNs due to the potential need to learn high frequencies or complex structures in the a priori unknown target function [99]. However, this inevitably increases training costs. Furthermore, oversampling on a high loss region may adversely slow convergence since it can be harder to learn such local complex gradient behaviour—a cause of stiffness during training [56, 57, 42, 43]. On the other hand, an overly small batch size will result in noisy (big jump in loss values) and inconsistent local gradient signals [100], thereby causing issues during gradient descent.
Figure 6: Example of learning PINN for simulating a distribution of tasks with a priori unknown input conditions in a what-if analysis: projectiles of ball with different physics properties and initial conditions (launch angle and launch speed).
III-A4 Current Research Directions
Since the seminal work of Raissi et al. [31], the architecture design and training methods for PINNs have continued to evolve. Recent advancements often combine multiple learning strategies to tackle the aforementioned PINN optimization issues [101, 102, 103]. Curriculum learning improves convergence by initiating PINNs training on a better-conditioned, simpler version of the original problem and then systematically advancing to target problem [39, 41]. Physics-informed learning often leads to premature convergence when dealing with challenging problems, however, the same PDE equation with an altered parameters or IC/BCs may become a simpler problem. In this scenario, one can start training the PINN model with the easier problem and then gradually vary the problem during training until reaching the target problem. This approach assumes that after each transition step, the training will bring the model close enough to the optimal solution of the next (harder) problem, providing a good starting point for the next round of training. Hence, curriculum learning can be viewed as an initialization process to guide the model toward the desired solution utilizing a priori understanding of how the system behaves as its problem parameters change. Adaptive loss-balancing strategy can also be considered as a form of curriculum learning, where the training is initiated with the loss weights corresponding to a simpler optimization problem. It involves assigning appropriate weights to balance different terms in the PINN loss functions [104, 86, 83, 50, 82, 84, 85], and is regarded as an important technique to improve convergence.
Recent studies have proposed enhancements to PINN learning capability by leveraging network architectures that incorporate Fourier feature embeddings, skip connections, and random weight factorization [101, 102, 103]. PINNs with skip connections and periodic activation functions retain a greater proportion of gradients during backpropagation from output to input, helping to mitigate spectral bias caused by activation saturation [105, 89, 91]. Adaptive techniques that adjust the slope of activation functions via trainable multiplicative factors have also been introduced to further control saturation and enhance training efficiency [106, 107]. These methods enable PINNs to embed inductive biases—such as high-frequency behavior characteristic of certain PDEs—directly into the network architecture. Additionally, physics-informed cell representations have been developed to address spectral bias away from the input origin by disentangling network parameters from the input coordinates [108]. Convolutional architectures also achieve this parameter-input decoupling, allowing for larger and more expressive networks that may scale more effectively to complex, large-scale problems [61].
Some studies have proposed using numerical differentiation (ND)-based methods for the differential operators required by the physics-informed loss, which is much cheaper to compute as compared to automatic differentiation-evaluated loss [56, 57, 62]. More importantly, ND are more robust to sample density and potentially advantageous in the context of reducing sample requirement in establishing a strong correlation between physics-informed loss and model accuracy. This is because ND methods, by definition, use Taylor series expansion (TSE) to construct a low-order local approximation for the derivative terms. They can provide structural biases and discourage extreme gradients to reduce overfitting and make PINNs easier to learn the approximate solution [56]. They can also better connect neighboring samples and facilitate the propagation of information from the IC/BC to interior sample points, thereby reducing any tendency to overfit especially near boundaries [57].
@@ III-B Challenges in Learning Generalizable PINN Model
Given both simulation (hypothesis evaluation) and inference capabilities, PINNs are versatile and can be applied to various applications common in science such as model-predictive-control, scenario planning, and design exploration [109, 110, 111, 112]; see the example in Fig. 6. The commonality to these applications is the need for multiple accurate and physics-compliant predictions across a series of related problems. These repeated evaluations usually differ by modifications to the physics scenario such as changes in geometric shape, PDE parameter, IC and BC.
@@ III-B1 How Generalizable Is A PINN Model?
Most parameterized PINN models (ref. to Subsection II-C; they are designed to learn a parametric family of physics scenarios) make predictions on new scenarios unseen during training through interpolation. Contrary to its promise, the interpolation is usually physics-agnostic, with similar negative implications for generalizability as the predictions from typical DNNs [55]. Given the difficulties in optimizing a PINN model for a single physics scenario, the cost of performing physics-informed learning across a set of scenarios can be prohibitive [113]. The effectiveness of these parameterized PINNs for prediction of new scenarios then depends on the complexity of the problem. Moreover, the problem’s parameterization needs to be determined at the learning stage, which reduces the flexibility of the parameterized PINN model when applied to new physics scenarios.
@@ III-B2 How to Train A Generalizable PINN Model in The Data-Scarce Regime?
IV Evo-PINN: Physics-Informed Learning with Evolutionary Algorithms
@@
Despite their theoretical advantage in diverse scientific domains, Section III highlighted several practical challenges in today’s PINN learning and optimization. The cost of optimization, navigating spectral bias and local minima traps, has even seeded a sense of pessimism among computational scientists about the utility of these techniques over standard numerical methods [124, 125]. In this section, we present our hypothesis that many algorithmic ideas developed in the computational intelligence community may in fact be uniquely well positioned to tackle the issues plaguing PINNs. One exciting research direction is to achieve physics-informed learning by synergizing evolutionary algorithms with local gradient descent signals, thereby leveraging the best of both worlds to quickly arrive at physically accurate and generalizable PINN models. In what follows, we examine this and other promising algorithmic directions that we believe warrant sustained investigation. These directions highlight key opportunities for the evolutionary optimization of PINNs (Evo-PINNs) to advance scientific machine learning.
IV-A Exploration-Oriented Evolutionary Algorithms for Tackling PINN Optimization Challenges
IV-A1 Neuroevolution Can Avoid Local Minima Traps
@@
In looking at a PINN’s unique optimization challenges, the need for new algorithms to overcome local minima traps caused by flat outputs at initialization, or by the attractiveness of unstable fixed point solutions to dynamical systems, becomes apparent. Gradient descent is a point-based optimization method that is highly sensitive to the issues of local minima or learning bias magnified by PINN loss functions. In contrast, in silico evolutionary computation offers a class of highly parallelizable approaches for guiding populations of candidate solutions towards globally-optimal regions of a search space via randomized mechanisms inspired by biological evolution [126]. We broadly refer to any such algorithmic realization as an evolutionary algorithm (EA) and its application to the training of neural networks as neuroevolution.
@@
The operators of EAs are typically based on principles of natural selection. They do not require derivative computations of the loss functions being optimized and are therefore gradient-free. Importantly, as population-based methods, the implicit parallelism of EAs implies that they have lesser tendency to get trapped in a single local minimum [127, 52]. What’s more, since their search is typically invariant to order-preserving transformations of the loss function [128], EAs are less susceptible to flat output functions with small (vanishing) gradients. For these reasons, they are deemed as potential alternatives to SGD, particularly in the context of scientific machine learning where the goal is to converge to physically precise solutions represented by global optima with zero PINN loss [51]. In addition to its global search capacity, an evolved population naturally provides an ensemble of neural networks that affords the possibility of accurate predictions with robust uncertainty quantification—as was shown through the application of a swarm-based evolutionary procedure to PINN optimization [129, 130]. The ability to provide uncertainty quantification and probabilistic forecasts is crucial for many scientific modelling and control problems [131].
IV-A2 Pareto Optimization Balances Physics and Data Losses
@@
The existence of multiple loss terms—see Eqs. (2) and (3)—makes PINN training intrinsically a multi-objective optimization problem. Since the terms $\mathcal{L}_{data},\mathcal{L}_{pde},\mathcal{L}_{ic}$ and $\mathcal{L}_{bc}$ represent different physical quantities, their magnitudes may be incommensurable. In this case, the role of the scalarizing weights is not only to express preferences between the losses, but also to compensate for differences in their magnitudes. It is worth noting that a perfect physics solution would, in theory, simultaneously satisfy all PINN loss terms. However, in practice, the expressive capacity of a neural network is always limited by its (necessarily) finite size. As argued in [97, 132], capacity limitations naturally give rise to conflicts and trade-offs between loss terms. This, in conjunction with the extreme ruggedness of the physics-informed learning objective, implies that a neural network, despite being a theoretically universal function approximator, is not guaranteed to find a solution that simultaneously minimizes all the terms [49]. On the other hand, when data-driven terms are included, such as observational data or empirical measurements, true contradictions may arise. This can occur when the data is noisy, incomplete, or governed by partially unknown physics that differs from the modeled equations. In such cases, the multi-objective optimization must reconcile inherently conflicting objectives between data fitting and physics adherence. Identifying an appropriate aggregation of objectives to arrive at good (physically satisfactory) trade-off solution is also non-trivial [133] Moreover, the shape of the Pareto front formed by the multiple objectives is not known beforehand [134], with simple weighted-sum aggregations being insufficient to support all parts of non-convex fronts [135].
@@
IV-A3 Memetic Algorithms Can Navigate Complex PINN Landscapes
There is a clear theoretical rationale supporting Evo-PINNs, yet there are only a handful of related works in the literature that explore such ideas. A barrier to the widespread use of EAs has been their high sample complexity, a consequence of updating and evaluating populations of solutions over multiple iterations. It is useful to consider hybridizing global evolutionary search with local gradient signals—thus inheriting the merits of both approaches. While the former can serve as an efficient basin-hopper for exploring rugged PINN landscapes, the latter facilitates rapid convergence to stationary points by exploiting local derivative information of loss functions.
@@ IV-A3 Memetic Algorithms Can Navigate Complex PINN Landscapes
The crafting of synergistic optimization frameworks to meld gradient-based and gradient-free learning poses fundamental challenges, including determining interaction mechanisms that best harness the strengths of the two approaches. Key questions to address include balancing the intensity of SGD versus EAs, ascertaining when gradient descent should take over from evolution, or selecting an appropriate subset of the evolving population for fine-tuning with gradient descent (as doing so for the entire population would be too expensive). Notably, the subfield of memetic computation in computational intelligence—where ‘memes’ are loosely defined as computationally-encoded units of information (e.g., local gradient signals) combined with global evolutionary search for effective problem-solving—offers a wealth of theories and methodologies that directly tackle these issues [147, 148].
IV-A4 Implementation Considerations and Hardware Acceleration
Employing EAs to directly optimize the neural network parameters introduces two primary challenges. First, EAs typically involve numerous hyperparameters (e.g., mutation rates, crossover rates, and population size) that require careful tuning. Fortunately, a substantial body of research has explored adaptive strategies for EA hyperparameter control [149], enabling EAs to better accommodate optimization problems of varying scales and characteristics. Among these, the Covariance Matrix Adaptation Evolution Strategy (CMA-ES) [150] stands out as a representative algorithm, widely applied across machine learning domains. It has also been implemented in several EA-based optimization frameworks, such as EvoJAX [151], which facilitate efficient deployment in machine learning tasks. These integrated and adaptive EA implementations help reduce the burden of manual hyperparameter tuning in neuroevolution applications. Second, the large-scale nature of neural network parameter spaces presents a significant challenge, as EAs often exhibit reduced efficiency in high-dimensional optimization problems. Nevertheless, as highlighted in [152], this does not imply that larger neural networks necessarily underperform smaller ones when optimized using EAs. In fact, the global search capability of EAs and their robustness to local minima can possibly enable even smaller networks to achieve competitive predictive performance. Furthermore, several advanced strategies—such as sparse optimization [153] and cooperative coevolution [154]—have been proposed to improve the scalability of EAs in large-scale settings. These methods offer valuable insights and serve as potential tools for addressing high-dimensional optimization challenges in the context of Evo-PINNs.
@@ IV-A4 Implementation Considerations and Hardware Acceleration
IV-B Neural Architecture Search for Inducing Physics Biases
The choice of architectures and hyperparameters for PINNs has often been guided by experiential priors drawn from standard deep learning. However, a growing body of literature shows that accurate physics-informed learning may call for bespoke architectures. For example, contrary to popular belief, it was found that more hidden layers do not necessarily mean better performance and can sometimes be detrimental. An automated neural architecture search via a bilevel programming formulation revealed that, for certain governing differential equations, a shallow neural network with many neurons was more suitable for PINNs [156]. Similar intriguing insights regarding the choice of activation function have also been uncovered. While squashing functions (e.g., tanh and sigmoid) and variants of the rectified linear unit are the mainstay in the deep learning literature, the use of Fourier features (i.e., sine and cosine) has gained prominence in physics tasks [89, 91], specifically for their ability to overcome spectral bias and capture periodic or high frequency patterns (e.g., shock fronts) common in various physics applications [157, 158, 159].
@@ IV-B Neural Architecture Search for Inducing Physics Biases
Figure 7: Results extracted from [51] where empirical comparisons between SGD and evolution strategies are provided. A PINN simulating projectile motion under gravitational forces on Earth, Mars, or the moon is optimized by transfer natural evolution strategies—dubbed tNES. The convergence trends from SGD, xNES, and tNES for selected scenarios are displayed (the bold lines indicate the mean convergence path of multiple independent runs and the shaded areas indicate their 10th-90th interpercentile ranges). Individual solutions given by multiple SGD and tNES (transfer from a relevant prior, i.e., source distribution obtained by solving a different gravity $g$ or launch angle $a_{0}$) optimization runs for one particular scenario ($g$=1.6, $a_{0}$=45) are plotted and compared against the ground truth.
IV-C Transfer Optimization for Accelerated PINN Training
IV-C1 Learning from Related Tasks
@@ IV-C2 Adaptive Transfer
IV-D Evolutionary Meta-Learning for Generalizable PINNs
IV-D1 Meta-Learning in PINNs
Meta-learning algorithms provide a systematic framework for reusing knowledge from a distribution of learning tasks. This allows a meta-learned PINN model to quickly adapt to new physics-informed learning tasks with minimal training. Crucially, meta-learning offers a pathway towards a generalizable PINN over a broad class of physics scenarios, including different PDE families and geometries, as highlighted in Subsection III-B2. For example, the MAML and Reptile algorithms have been recently explored in the context of PINN [118, 119, 117]. They seek to meta-learn an optimal model initialization in the space of network weights from which the network can rapidly converge to a solution that minimizes the physics-informed loss of new tasks. This is achieved by a task-specific fine-tuning procedure during the meta-learning stage through an inner-outer loop training framework. This problem can be mathematically posed as a two-stage or bi-level optimization. The network weights are trained explicitly (in the outer loop) by optimizing the model performance over a batch of physics-informed learning tasks after one or a few gradient descent updates (at the inner loop) from the learned initialization.
@@ IV-D2 Neural Baldwinism
Figure 9: Schematic to illustrate new tasks arising from family of PDEs problem: S1 change to new PDE ($\gamma$ and $\delta$ parameters), source term $q$, and IC profile for $t\in[0,2]$ (same time domain as train tasks), and S2 projection to longer time domain $t\in[0,4]$. The solution for new PDE tasks (they are visually indistinguishable from the ground truth) can be quickly obtained in the order of milli-seconds by the Baldwinian-PINN [55]. In [55], CMA-ES was employed to evolve the Baldwinian-PINN models, demonstrating effective convergence in both physics residuals and MSE (108 training tasks). Superior performance is achieved with CMA-ES using initial standard deviations of 5 and 10. The mean MSE across 87 test tasks is below 5e-5.
Figure 10: Empirical comparison between results reported in Baldwinian-PINN [55] and recent meta-learning PINN studies. For linear problems, the results for NRPINN and Hyper-LR-PINN are extracted from [117] and [119], respectively. For nonlinear problems, the results for other meta-learning PINN models / methods (inc. MAML) are extracted from [118]. On nonlinear test tasks, evolved Baldwinian-PINNs achieve up to 10x lower relative norm error, with prediction times under 2 seconds—over 100x faster than other meta-learning methods. For instance, in diffusion-reaction experiments, Baldwinian-PINN yields 70x better accuracy with 700x less computation time.
@@
Recently, a meta-learning PINN framework based on the evolutionary principle of Baldwinism—dubbed Baldwinian PINN [55]—has been proposed to address the training challenges faced in conventional gradient-based meta-learning algorithms such as MAML. The Baldwin effect in nature explains the development of phenotypic plasticity among a population’s individuals [188], whereby their young are “born ready” with strong learning biases to perform a wide range of tasks. In order to pre-wire such learning ability into the initial connection strengths of a neural network, enabling it to be performant over a distribution of physics-informed learning tasks, an algorithmic realization of neural Baldwinism is examined in the context of PINNs. The algorithm tackles the two-stage optimization problem in meta-learning by coupling hidden layer evolution, for better generalization across tasks, with lifetime-learning at the output layer to specialize to particular tasks. Task-specific learning is limited to the output layer only [189, 190, 191], allowing the model to solve new linear PDE problems with a one-step pseudoinverse operation (or a few pseudoinverse steps for nonlinear PDE problems). The pseudoinverse operation has a closed form and is therefore fast. Unlike gradient-based MAML, the global evolutionary search in Baldwinian-PINNs is better equipped to handle the two-stage/bi-level problem formulation where exact gradients may be difficult to compute. Being a gradient-free algorithm, it does not require unrolling and differentiating through the computation graph, making Baldwinian-PINN an efficient method that scales well with the number of GPUs available. Fig. 8 illustrates the contrast between MAML and Baldwinian evolution for PINN [^2] Python source codes for Baldwinian-PINN are available at https://github.com/chiuph/Baldwinian-PINN..
@@
Experiments in [55] clearly illustrate how Baldwinian-PINN can consistently solve new physics scenarios within a diverse task distribution with orders of magnitude improvement in speed and accuracy compared to a single PINN model. These PDEs include both linear equations such as the linear convection-diffusion equation, Helmholtz equation and Poisson equation, and non-linear equations such as the Burger’s equation, Allen-Cahn equation and diffusion-reaction equation. Examples of the effective generalization of a Baldwinian-PINN to an entire PDE family with diverse output patterns is demonstrated in Fig. 9, on 2 task scenarios: S1 shows successful learning of solution for unseen set of PDEs, which include changes to PDE parameters, source term, and IC profile; while S2 shows effective extrapolation of solution to a longer time domain. Baldwinian-PINNs have clear benefits for the continual modeling of dynamical systems through their versatility in handling ICs and ability to rapidly model and stitch time windows together with minimal error. Crucially, Baldwinian evolution does not require explicit parameterization, and evolved Baldwinian-PINN can flexibly generalize to new ICs, BCs, and PDE source terms, a stark contrast to the need for existing parameterized PINNs to have pre-defined PDE, BC or IC parameters. Overall, the outstanding results in [55] compared to other meta-learning methods (Fig. 10) strongly motivate similar approaches in the spirit of Baldwinian-PINN.